Given that P = (8, 8) and Q = (9, 17), find the component form and magnitude of the vector PQ.


1
Expert's answer
2021-06-07T19:07:25-0400

"Given

\\P=(8,8)\\Q=(9,17)"


So, "PQ=(9-8,17-8)=(1,9)"


And now we can find the magnitude of our vector:

"|PQ|=\sqrt{(1)^2+(9)^2}=\sqrt{1+81}=\sqrt{82}"


Answer:


PQ = (1,9) and |PQ| = "\sqrt{82}"